# Learning on LLM Output Signatures for gray-box LLM Behavior Analysis

### v1

@@ Acknowledgments
Appendix A Appendix
Appendix B Proofs
@@
The finite disjoint union of compact subsets correspond to all possible sequence lengths ($N^{\prime}\in N_{\text{max}}$) and vocabulary sizes ($V^{\prime}\in V_{\text{max}}$). Below we provide the proof for Lemma B.3.
@@
Appendix C Extended Experimental Section
C.1 Experimental Details
Our experiments were conducted using the PyTorch [36] framework, using NVIDIA L40 GPUs. We use a fixed batch size of 64 for all the tasks and datasets, and a fixed value of 8 heads (except for the Movies[35] dataset) in our light-weight transformer encoder for LOS-Net. Hyperparameter tuning was performed utilizing the Weight and Biases framework [5] – see Table 4.
C.2 HyperParameters
@@ C.3 Optimizers and Schedulers
C.4 Our Baselines and Rank Encoding
@@
C.5 Dataset Description
C.5.1 Datasets for Hallucination Detection
@@
    HotpotQA [52]: This dataset is specifically designed for multi-hop question answering and includes diverse questions that require reasoning across multiple pieces of information. Each entry comprises supporting Wikipedia documents that aid in answering the questions. For our analysis, we utilized the “without context” setting, where questions are posed directly. This setup demands both factual knowledge and reasoning skills to generate accurate answers.
@@
    Movies [35]: To evaluate generalization in scenarios regarding movies involving factual inaccuracies (i.e., hallucinations), we employed the dataset introduced by [35].
@@
    IMDB [30]: This dataset contains movie reviews designed for sentiment classification tasks. Following the approach outlined in [35], we applied a one-shot prompt to guide the large language model (LLM) in using the predefined sentiment labels effectively.
@@
C.5.2 Datasets for Data Contamination Detection
BookMIA. The original BookMIA data have been obtained from the Hugging Face dataset swj0419/BookMIA [^8] https://huggingface.co/datasets/swj0419/BookMIA., accessed via the Hugging Face python datasets API. The dataset totals $9,870$ excerpts from a total of $100$ books, of which $50$ are labeled as members (positives) and $50$ are labeled as non-members (negatives).
@@ C.5.2 Datasets for Data Contamination Detection
In the case of the reference-based baseline, we consider the smallest-sized available counterparts for the respectively attacked LLMs, namely: Pythia 70M for Pythia models and Llama-1 7B for Llama models. All LLMs are accessed through the Hugging Face python interface, specifically: EleutherAI/pythia-70m, EleutherAI/pythia-{6.9,12}b [^9] https://huggingface.co/EleutherAI/pythia-70m, https://huggingface.co/EleutherAI/pythia-6.9b, https://huggingface.co/EleutherAI/pythia-12b. and huggyllama/llama-{7,13,30}b [^10] https://huggingface.co/huggyllama/llama-7b, https://huggingface.co/huggyllama/llama-13b, https://huggingface.co/huggyllama/llama-30b..
WikiMIA. WikiMIA[41] is the first benchmark for pre-training data detection, comprising texts from Wikipedia events. The distinction between training and non-training data is determined by timestamps. WikiMIA organizes data into splits based on sentence length, enabling fine-grained evaluation. It also considers two settings: original and paraphrased. The original setting evaluates the detection of verbatim training texts, while the paraphrased setting, where training texts are rewritten using ChatGPT, assesses detection on paraphrased inputs. In this paper, we consider the original (non-paraphrased) split and focus on the 32 and 64 split sizes, as they contain the largest number of samples, approximately 750 and 550, respectively.
C.6 Ablation Study
@@
The Role of the TDS ($\mathbf{X}$). As a case study, we examine both the DCD task on the BookMIA dataset and the HD task across the three datasets: HotpotQA, IMDB, and Movies. Our analysis involves six LLMs. To examine to role of the TDS, we benchmark LOS-Net against our two proposed baselines, which focus on the ATP, namely, ATP+R-Transf. and ATP+R-MLP.
The results, summarized in Figure 7, consistently demonstrate that the best-performing model is LOS-Net. In many cases, LOS-Net outperforms the alternatives by a significant margin, indicating that the information encoded in the TDS ($\mathbf{X}$) is crucial for both tasks.
When comparing the two ATP-based baselines, ATP+R-Transf. and ATP+R-MLP, we find that in 8 out of 10 cases, ATP+R-Transf. achieves better performance. This suggests that treating ATP ($\mathbf{p}$) as sequential data - the more natural approach - is more effective than using an MLP.
The hyperparameter $K$. To evaluate the impact of the hyperparameter $K$ introduced in Equation 1, we conduct a comprehensive case study focusing on the task of HD. This analysis is performed across two LLMs, Llama3-8b-instruct and Mistral-7b-instruct, as well as two diverse datasets, HotpotQA and IMDB.
@@
Figure 6: Ablation study analyzing the effect of the hyperparameter $K$ introduced in Equation 1. The gap between the highest and lowest bin is indicated on the top right of each histogram.
It is evident that the hyperparameter $K$ has little impact on the final results. Specifically, three out of the four gaps are less than 2 AUC, and the largest gap is below 5 AUC. Surprisingly, using the full vocabulary (the rightmost bin in all histograms) does not show any clear performance improvement but significantly slows down training compared to $K\leq 1000$. Thus, as an optimal balance between training time and performance, we select $K=1000$ for all the experiments unless otherwise specified.
Figure 7: Ablation study evaluating the role of the TDS ($\mathbf{X}$) and the ATP ($\mathbf{p}$). Results are shown for the four LLMs: L-3-8b, Pythia-12b, Mis-7b, and LLaMa-13b.
C.7 Extended Results for Transferability
Here we present the complementary results for Figures 4 and 5, for the case where we trained from scratch. The results are presented in Figures 8 and 9.
Figure 8: Training from scratch performance (AUC) in the Cross-LLM setting. Y-axis: source LLMs (train), X-axis: target LLMs (test).
Figure 9: Training from scratch performance (AUC) in the cross-dataset setting. Y-axis: train datasets, X-axis: test datasets.
Appendix D Additional Tasks Background
@@
For HD, ground-truth labels can be collected by providing the target LLM with inputs prompting for completion or question answering on known facts and/or reasoning tasks. Hallucinations or error annotations are derived by comparing the consistency of the model’s response with known, factually true, or logically correct answers. For further details, refer to Section C.5.1.

### v2

@@ Acknowledgments
Appendix A Proofs
@@
The finite disjoint union of compact subsets correspond to all possible sequence lengths ($N^{\prime}\in N_{\text{max}}$) and vocabulary sizes ($V^{\prime}\in V_{\text{max}}$). Below we provide the proof for Lemma A.3.
@@
Appendix B Extended Experimental Section
B.1 Experimental Details
Our experiments were conducted using the PyTorch [36] framework (License: BSD), using a single NVIDIA L-40 GPU for all experiments regarding LOS-Net. We use a fixed batch size of 64 for all the tasks and datasets, and a fixed value of 8 heads (except for the Movies[35] dataset) in our light-weight transformer encoder for LOS-Net. Hyperparameter tuning was performed utilizing the Weight and Biases framework [5] – see Table 3.
B.2 HyperParameters
@@ B.3 Optimizers and Schedulers
B.4 Our Baselines and Rank Encoding
@@
B.5 Dataset Description
B.5.1 Datasets for Hallucination Detection
@@
    HotpotQA [54] (License: CC-BY-SA-4.0): This dataset is specifically designed for multi-hop question answering and includes diverse questions that require reasoning across multiple pieces of information. Each entry comprises supporting Wikipedia documents that aid in answering the questions. For our analysis, we utilized the “without context” setting, where questions are posed directly. This setup demands both factual knowledge and reasoning skills to generate accurate answers.
@@
    Movies [35] (License: MIT): This dataset checks for factual accuracy in scenarios regarding movies. LLMs are asked, in particular, who was the actor/actress playing a specific role in a movie of interest. This dataset contains 7857 test samples.
@@
    IMDB (originally released with no known license by Maas et al. [30]): This dataset contains movie reviews designed for sentiment classification tasks. Following the approach outlined in [35], we applied a one-shot prompt to guide the large language model (LLM) in using the predefined sentiment labels effectively.
@@
LLMs. We consider the following LLMs for our experiments on HD:
  1. 1.
    Mistral-7b-instruct-v0.2 [21] (License: Apache-2.0). Referred to as Mis-7b in the main text and accessed through the Hugging Face interface at https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3.
  2. 2.
    Llama-3-8b-Instruct [46] (License: Llama-3777https://huggingface.co/meta-llama/Meta-Llama-3-8B/blob/main/LICENSE). Referred to as L-3-8b in the main text and accessed through the Hugging Face interface at https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct3.
  3. 3.
    Qwen-2.5-7b-Instruct (License: Apache-2.0): Referred to as Q-2.5-7b in the main text and accessed through the Hugging Face interface at https://huggingface.co/Qwen/Qwen2.5-7B-Instruct.
B.5.2 Datasets for Data Contamination Detection
BookMIA. [41] The original BookMIA data have been obtained from the Hugging Face dataset swj0419/BookMIA [^8] https://huggingface.co/datasets/swj0419/BookMIA., accessed via the Hugging Face python datasets API (License: MIT). The dataset totals $9,870$ excerpts from a total of $100$ books, of which $50$ are labeled as members (positives) and $50$ are labeled as non-members (negatives).
@@ B.5.2 Datasets for Data Contamination Detection
In the case of the reference-based baseline, we consider the smallest-sized available counterparts for the respectively attacked LLMs, namely: Pythia 70M for Pythia models and Llama-1 7B for Llama models. All LLMs are accessed through the Hugging Face python interface, specifically: EleutherAI/pythia-70m, EleutherAI/pythia-{6.9,12}b [^9] https://huggingface.co/EleutherAI/pythia-70m (License: Apache-2.0), https://huggingface.co/EleutherAI/pythia-6.9b, https://huggingface.co/EleutherAI/pythia-12b. and huggyllama/llama-{7,13,30}b [^10] https://huggingface.co/huggyllama/llama-7b, https://huggingface.co/huggyllama/llama-13b, https://huggingface.co/huggyllama/llama-30b. (License: Llama [^11] https://huggingface.co/huggyllama/llama-13b/blob/main/LICENSE, https://huggingface.co/huggyllama/llama-30b/blob/main/LICENSE).
WikiMIA. WikiMIA[41] (License: MIT) is the first benchmark for pre-training data detection, comprising texts from Wikipedia events. The distinction between training and non-training data is determined by timestamps. WikiMIA organizes data into splits based on sentence length, enabling fine-grained evaluation. It also considers two settings: original and paraphrased. The original setting evaluates the detection of verbatim training texts, while the paraphrased setting, where training texts are rewritten using ChatGPT, assesses detection on paraphrased inputs. In this paper, we consider the original (non-paraphrased) split and focus on the 32 and 64 split sizes, as they contain the largest number of samples, approximately 750 and 550, respectively. On top of the LLMs attacked in BookMIA, here we also attack Mamba-1.4b (License: Apache-2.0), accessed via the Hugging Face interface (https://huggingface.co/state-spaces/mamba-1.4b).
B.6 Ablation Study
@@
Figure 6: Ablation study evaluating the role of the TDS ($\mathbf{X}$) and the ATP ($\mathbf{p}$) on our HD setups, including datasets HotpotQA, IMDB, Movies, and LLMs L-3-8b, Mis-7b, Q-2.5-7b.
Figure 7: Ablation study evaluating the role of the TDS ($\mathbf{X}$) and the ATP ($\mathbf{p}$) on BookMIA for Pythia and Llama-1 LLMs.
The Role of the TDS ($\mathbf{X}$). As a case study, we examine both the DCD task on the BookMIA dataset and the HD task across the three datasets: HotpotQA, IMDB, and Movies. Figures 6 and 7 report a close-up comparison between LOS-Net and our two proposed baselines, which explicitly neglect the TDS, namely, ATP+R-Transf. and ATP+R-MLP. These plots consistently show how the best-performing model is LOS-Net. In many cases, LOS-Net outperforms the alternatives by a significant margin, indicating that the information encoded in the TDS ($\mathbf{X}$) is crucial for both tasks. Regarding the two ATP-based baselines, we report that ATP+R-Transf. obtains better performance than ATP+R-MLP in 8 out of 12 cases, but these improvements do not seem to follow a clear pattern across LLMs and datasets. The only exception is BookMIA, on which the former architecture outperformed the latter across all the four attacked LLMs.
Figure 8: Ablation study analyzing the effect of the hyperparameter $K$ introduced in Equation 1.
The hyperparameter $K$. To evaluate the impact of the hyperparameter $K$ introduced in Equation 1, we conduct a comprehensive case study focusing on the task of HD.
We experiment with various values of $K$, specifically $K\in\{10,50,100,500,1000\}$, and trained the same selected model whose results are reported in Table 1 for $K=1000$. The corresponding Test AUCs are presented in Figure 8.
From the reported bar plots, we do observe that performances either weakly increase with $K$ (see, e.g., Movies for Q-2.5-7b or HotpotQA on L-3-8b), or stay approximately constant (see, e.g., IMDB on Mis-7b). In any case, the performance difference w.r.t. our default setting $K=1000$ remains contained. This is a valuable feature, as it unlocks the effective application of LOS-Net even on non fully open LLMs such as the most recent models released by OpenAI [^12] At the time of writing, OpenAI’s API only gives access to a maximum of $20$ top scoring logprobs (https://platform.openai.com/docs/api-reference/completions/create, accessed May 2025..
Table 4: Test AUC scores for HD on Qwen-2.5-7b-Instruct (Q-2.5-7b). The best-performing method is in bold, and the second best is underlined.
Method HotpotQA IMDB Movies
Q-2.5-7b
Logits-mean 66.2 74.8 71.3
Logits-min 59.8 72.1 42.1
Logits-max 60.4 60.7 65.1
Probas-mean 67.5 74.6 74.2
Probas-min 54.4 65.4 44.7
Probas-max 61.8 50.1 72.9
ATP+R-MLP 71.38
$\pm$ 0.28
84.69
$\pm$ 0.37
78.06
$\pm$ 0.45
ATP+R-Transf. 69.34
$\pm$ 2.04
87.73
$\pm$ 0.03
77.37
$\pm$ 3.13
LOS-Net 73.71
$\pm$ 1.21
88.19
$\pm$ 0.88
88.00
$\pm$ 0.39
B.7 Results For Hallucination Detection for Qwen-2.5-7b
Table 4 reports results on our three considered HD datasets over LLM Qwen-2.5-7b-Instruct (Q-2.5-7b) [53]. We can see LOS-Netoutperforms all non-learnable output-based baselines by large margin, as well as our learnable baselines ATP+R-Transf. and ATP+R-MLP.
B.8 Results On The WikiMIA Dataset
Dataset $\rightarrow$ WikiMIA - 32 WikiMIA - 64
LLM $\rightarrow$ P-6.9b L-13b L-30b M-1.4b P-6.9b L-13b L-30b M-1.4b
@@ B.8 Results On The WikiMIA Dataset
As shown in Table 5, LOS-Net consistently surpasses all baseline methods across all eight combinations of LLMs and datasets. Notably, for L-30b, our model achieves an AUC score that is more than 8 points higher than the best-performing baseline, MinK%++ for both datasets, demonstrating a substantial improvement. Similarly, for P-6.9b, our model maintains a steady advantage of approximately 5 AUC for both datasets, further underscoring its robustness. Overall, the second-best method is MinK%++, followed by MinK%, consistent with the findings of [58].
B.9 LOS-Net Run-Time
In Table 6, we report the wall-clock training times (for the best selected model based on the held-out validation set) and single-example detection times for LOS-Net  for all experiments presented in this paper.
Table 6: Comparison of training and detection times of our model LOS-Net  across all the DC settings explored in our paper, as well as HD settings for Mis-7b and L-3-8b. All are measured on a single NVIDIA L-40 GPU.
Task Target LLM Dataset Training Time [h = hours, m = minutes, s = seconds] Detection Time (Mean $\pm$ Std) [seconds] HD Mis-7b HotpotQA 9m 19s $3.32\times 10^{-5}\pm 1.20\times 10^{-5}$ s IMDB 16m 8s $4.05\times 10^{-5}\pm 1.83\times 10^{-5}$ s Movies 17m 50s $1.95\times 10^{-5}\pm 7.24\times 10^{-6}$ s L-3-8b HotpotQA 6m 39s $2.38\times 10^{-5}\pm 7.18\times 10^{-6}$ s IMDB 4m 23s $3.37\times 10^{-5}\pm 1.53\times 10^{-5}$ s Movies 11m 34s $3.05\times 10^{-5}\pm 1.21\times 10^{-5}$ s DCD L-13b WikiMIA-32 33m 6s $4.13\times 10^{-5}\pm 1.67\times 10^{-6}$ s WikiMIA-64 2m 7s $2.67\times 10^{-5}\pm 1.12\times 10^{-5}$ s BookMIA 7m 32s $3.67\times 10^{-5}\pm 8.65\times 10^{-6}$ s L-30b WikiMIA-32 28m 40s $4.05\times 10^{-5}\pm 3.10\times 10^{-6}$ s WikiMIA-64 5m 8s $4.96\times 10^{-5}\pm 2.54\times 10^{-5}$ s BookMIA 16m 38s $3.94\times 10^{-5}\pm 1.42\times 10^{-5}$ s P-6.9 WikiMIA-32 24m 55s $2.91\times 10^{-5}\pm 4.41\times 10^{-6}$ s WikiMIA-64 26m 13s $3.18\times 10^{-5}\pm 1.56\times 10^{-5}$ s BookMIA 18m 23s $2.86\times 10^{-5}\pm 6.16\times 10^{-6}$ s P-12b BookMIA 19m 49s $4.07\times 10^{-5}\pm 4.89\times 10^{-6}$ s M-1.4b WikiMIA-32 1h 6m 18s $3.87\times 10^{-5}\pm 1.27\times 10^{-6}$ s WikiMIA-64 1h 16m 51s $3.12\times 10^{-5}\pm 1.42\times 10^{-5}$ s
Appendix C Additional Tasks Background
@@
For HD, ground-truth labels can be collected by providing the target LLM with inputs prompting for completion or question answering on known facts and/or reasoning tasks. Hallucinations or error annotations are derived by comparing the consistency of the model’s response with known, factually true, or logically correct answers. For further details, refer to Section B.5.1.
Appendix D Broader Impact
By improving the detection of data contamination and hallucinations in Large Language Models, our contribution could foster the responsible development and use of Generative AI by enhancing transparency and trustworthiness. Applications to detecting machine-generated content could also support reducing the spread of AI-fabricated misinformation. We also note, however, that our work sheds light on predictive information contained in the output next-token probability distributions. This could lead malicious actors to develop more sophisticated defense mechanisms to MIAs or lead companies to potentially limit access to LLM outputs, thus hindering advancements in open research.
Appendix E LOS-Net Visualization
In Figure 9 we provide a visualization of our architecture, LOS-Net .
Figure 9: A visualization of LOS-Net .